# Long Chain-of-Thought for Reasoning Large Language Models

### v1

@@ 7.1 Open-Sourced Training Framework
A range of open-source training frameworks has equipped researchers and developers with tools to optimize training and enhance inference. Each framework is built on distinct design principles and features. Early frameworks like SimpleRL [712] and DeepScaler [359] quickly replicated R1’s technology stack. Others, such as X-R1 [519] and TinyZero [406], emphasize delivering an intuitive “Aha moment” experience for under \$50. Open-Reasoner-Zero [183] replicated the DeepSeek-R1-zero training scheme with a 32B model and achieved a similar performance. Additionally, LLM Reasoner [161] provides tools to help researchers adapt strategies for External Exploration. Frameworks such as OpenR [557], OpenRLHF [182], OpenR1 [507], and Logic-RL [622] have enhanced the replication of Long CoT in deep reinforcement learning for text modalities. R1-V [62], R1-Multimodal-Journey [465], VL-Thinking [57], VLM-R1 [467], Open-R1-Multimodal [255], and Video-R1 [518] have extended the R1 framework to multimodal settings, enabling cross-modal R1-like reinforcement learning-based training. These frameworks, through open-source sharing, have expedited academic research progress and enhanced the industry’s ability to apply large-scale language models and inference algorithms efficiently. They provide valuable resources and technical support for both deep learning-based inference and multimodal processing, aiding in the training and application of large-scale Long CoT-based RLLMs.
@@ 8.1 Multimodal Long CoT
Recent discussions have focused on extending reasoning chains to multimodal contexts in the areas of Long CoT and multimodal reasoning [436, 373, 626, 613, 721, 706, 348, 169]. Zhang et al. [757] introduce multimodal chain-of-thought (MMCoT), while M3CoT [65] extends this with complex MMCoT, similar to Long CoT, and provides an evaluation benchmark. This work suggests that mimicking human Long CoT offers an effective solution [192, 163]. Multimodal Long CoT can be categorized into three main approaches: (1) Multimodal Long CoT Prompting: Eariler, Chen et al. [65] demonstrate that the basic description-then-reasoning prompt fails in Long CoT scenarios. To fill this gap, Li et al. [307] improve Vision RLLMs by enabling detailed, context-aware descriptions through an iterative self-refinement loop, allowing interactive reasoning for more accurate predictions without additional training. Dong et al. [105] incorporate multi-agent interaction during prompting, further scaling the reasoning length and achieving better accuracy. Furthermore, FaST [487] uses a switch adapter to select between Long CoT and direct answer modes, resulting in enhanced performance. (2) Multimodal Long CoT Imitation: Recent models such as LLaVA-CoT [633] and Virgo [112] employ data distillation to enable the imitation of Long CoT processes, addressing more complex problem-solving tasks [521]. Additionally, AtomThink [619] offers a Long CoT annotation engine that generates high-quality CoT annotations, mitigating the issue of insufficient visual mathematical data. Wei et al. [593] further extend Long CoT paradigms by incorporating more tokens during perception, improving geometric reasoning. (3) Reward Model-Based Multimodal Long CoT Exploration: Recent research employs reward or value models to enhance inference test-time scaling in both exploration and training phases. This includes model decoding [344, 42, 629] and training [619, 574, 718, 545], as well as the diffusion process [365], all contributing to improved visual reasoning and comprehension.
The primary challenges in multimodal Long CoT are: (1) Incorporating Multimodal Reasonings: Enabling RLLMs to assist reasoning by generating visual content holds promise for improving complex spatial reasoning tasks, particularly when logic cannot be easily conveyed through text alone [85, 276, 157]. (2) Extending Longer Reasoning Processes: While current models focus on imitating Long CoT, there remains a lack of exploration into how multimodal inference test-time scaling can be achieved through methods like RL or MCTS [605, 209], presenting an interesting avenue for future research.
@@ 8.3 Agentic & Embodied Long CoT
Researchers have expanded Long CoT in interactive environments, significantly improving success rates in automated exploration tasks [160, 767, 718]. Current research primarily focuses on two approaches: (1) Tree-based Search Augmentation Early work by Hao et al. [160], Koh et al. [249] introduce tree search techniques to enhance agent exploration. Hu et al. [185] further propose planning sampling strategies to accelerate tree search processes. Additionally, Light et al. [317] develop a method to gather high-quality interactive feedback through self-play simulations with MCTS and LLM-based reflection, which helps acquire high-level strategic skills and guide low-level execution. (2) Environmental Interactivity Improvement A key feature of Agentic Systems is their interaction with the environment, making the enhancement of this aspect a critical focus [160, 777, 244, 120]. Nie et al. [397] and Hu et al. [184] improve interactivity by incorporating memory history into the agent’s functions. (3) Multi-agent Cooperative Improvement Another key feature of agentic systems is that it can incorporate multiple agents to cooperative to solve a complex problem [796, 558, 427, 614, 794]. Christakopoulou et al. [90] introduce the Talker-Reasoner architecture, which separates the agent’s tasks into deep reasoning and rapid dialogue generation, providing a more effective interaction protocol. Lei et al. [270] introduce the Multi-Agent System for Conditional Mining (MACM) prompting method, which effectively addresses complex mathematical problems and exhibits robust generalization across diverse mathematical contexts.
@@ 8.4 Efficient Long CoT
The deep reasoning, exploration, and reflection of the Long CoT often lead to long outputs, which necessitate improved speedup techniques [662, 189, 134, 482, 212]. Consequently, optimizing reasoning for faster reasoning with maximum accuracy has become a significant challenge for Long CoT. Current research mainly focuses on two approaches: (1) Direct Compression and Shortening of Reasoning Chains: The most direct strategy is to consider direct compression and reducing the length of the reasoning chain while maintaining accuracy [86, 489, 20]. Specifically, a series of work [508, 357, 49, 368, 394] encourage the generation of shorter reasoning processes, minimizing redundancy and enhancing efficiency [17, 639, 27, 392, 570]. Additionally, researchers further introduce token budgets in prompts to control reasoning complexity, further improving efficiency [158, 711, 541, 211, 281, 5]. Building on these approaches, MARP [64] and DynaThink [404] allow LLMs to adapt reasoning speed based on task complexity, perplexity, or confidence, optimizing both efficiency and accuracy [147, 464, 799, 102, 98, 565, 235]. Moreover, Botta et al. [38] and Xia et al. [617] introduce a technique that enables LLMs to erase or skip some generated tokens, thereby compressing the reasoning length. More radically, Yu et al. [689] and Du et al. [109] propose distilling long reasoning paradigms into direct prediction models, reducing computational costs without sacrificing reasoning quality. (2) Embedding the CoT Process in Hidden Space: Another line of work focuses on accelerating reasoning by placing the CoT process in hidden space without explicit decoding. Specifically, Coconut [162], LaTRO [56], and SoftCoT [641] transfer reasoning into continuous latent space, promoting "continuous thinking" and enabling the model to maintain multiple alternative reasoning paths [732]. Similarly, Wang et al. [575] use "planning tokens" to enhance reasoning, performing the planning process in hidden space to save computational resources and improve inference performance.
@@ 8.5 Knowledge-Augmented Long CoT
The reasoning model significantly enhances reasoning capabilities, but it still lacks knowledge in specialized fields and timely new information [66, 117, 338]. Thus, enriching reasoning with additional knowledge presents a key challenge for Long CoT [60, 54]. Current research focuses primarily on two approaches: (1) Retrieval-Augmented Generation: Retrieval-Augmented Generation (RAG) techniques enhance LLMs by integrating dynamic knowledge retrieval and document refinement. Research has combined RAG with reasoning modules to improve performance on complex tasks [512, 298, 576, 150, 221, 226, 337, 609]. O1 Embedder [644] optimizes multi-task retrieval and reasoning through synthetic data training. Furthermore, Stream of Search (SoS) [127] CoRAG [564] boost search accuracy and addresses unresolved issues by incorporating more natural reflection and exploration in RAG. (2) Model Knowledge Injection: An alternative approach involves integrating additional knowledge during SFT or RL. Specifically, HuatuoGPT-o1 [60] utilize the R1-like paradigm to train LLMs by model-judged reward RL, which significantly improves the medical knowledge during reasoning [407]. Huang et al. [201] and Wang et al. [549] optimize for injecting medical knowledge in Long CoT scenarios by SFT, which also achieve great performance. Further, Jiang et al. [223] introduce MCTS to synthesize data, achieving superior performance. This model merges verifiable medical knowledge with reinforcement learning techniques to enhance performance in complex, medical task settings.
@@
8.6 Safety for Long CoT
Despite the significant performance improvements from Long CoT, Long CoT-augmented LLMs still face major safety challenges, particularly in generating unsafe outputs, such as misinformation and offensive content [786, 273, 785, 354, 18, 30, 29, 106, 241, 743]. Current research primarily addresses two key approaches: (1) Long CoT Attack Several studies show that Long CoT makes models more vulnerable to unexpected behavior [119] or unsafe outputs [254, 797, 638]. For instance, Arrieta et al. [19] identify that DeepSeek-R1 is prone to generating harmful content, including misinformation and offensive speech. Additionally, Kumar et al. [251] introduce the OverThink attack, which exploits false inference problems to induce overthinking in models, providing insights into potential defensive strategies. Further, Yao et al. [671] fool RLLMs chain of iterative chaos, for better jailbreaking. (2) Long CoT Safety Improvement Another major area of research focuses on enhancing safety [219, 793, 345] and reliability [450, 533] through prompting or training techniques. Shen et al. [469] present Heima, which optimizes inference efficiency and robustness. Gallego [125] proposes dynamic security prompts during inference, while Cheng et al. [84] address hallucinations by guiding reasoning with a tree search algorithm. Zhao et al. [763] introduce a self-reflection framework to identify biases, and Wang et al. [554] propose Safety Reasoning with Guidelines (SRG) to defend against out-of-distribution attacks. Finally, Parmar and Govindarajulu [415] combine reinforcement learning (RL) and supervised fine-tuning (SFT) in a hybrid training approach to reduce harmful outputs and enhance DeepSeek-R1’s safety.

### v3

@@ 7.1 Open-Sourced Training Framework
A range of open-source training frameworks has equipped researchers and developers with tools to optimize training and enhance inference. Each framework is built on distinct design principles and features. Early frameworks like SimpleRL [809] and DeepScaler [408] quickly replicated R1’s technology stack. Others, such as X-R1 [583] and TinyZero [456], emphasize delivering an intuitive “Aha moment” experience for under \$50. Open-Reasoner-Zero [209] replicated the DeepSeek-R1-zero training scheme with a 32B model and achieved a similar performance. Additionally, LLM Reasoner [186] provides tools to help researchers adapt strategies for External Exploration. Frameworks such as OpenR [624], OpenRLHF [208], OpenR1 [571], and Logic-RL [701] have enhanced the replication of Long CoT in deep reinforcement learning for text modalities. Further, DAPO [781] and VAPO [801] DAPO [781] and VAPO [801] enhance the efficiency of Long CoT RL training by incorporating more detailed and fine-grained training strategies. R1-V [74], R1-Multimodal-Journey [522], VL-Thinking [67], VLM-R1 [526], Open-R1-Multimodal [289], and Video-R1 [582] have extended the R1 framework to multimodal settings, enabling cross-modal R1-like reinforcement learning-based training. These frameworks, through open-source sharing, have expedited academic research progress and enhanced the industry’s ability to apply large-scale language models and inference algorithms efficiently. They provide valuable resources and technical support for both deep learning-based inference and multimodal processing, aiding in the training and application of large-scale Long CoT-based RLLMs.
@@ 8.1 Multimodal Long CoT
Recent discussions have focused on extending reasoning chains to multimodal contexts in the areas of Long CoT and multimodal reasoning [491, 423, 705, 689, 818, 802, 395, 194, 717, 419]. Zhang et al. [858] introduce multimodal chain-of-thought (MMCoT), while M3CoT [79] extends this with complex MMCoT, similar to Long CoT, and provides an evaluation benchmark. This work suggests that mimicking human Long CoT offers an effective solution [221, 188]. Multimodal Long CoT can be categorized into three main approaches: (1) Multimodal Long CoT Prompting: Eariler, Chen et al. [79] demonstrate that the basic description-then-reasoning prompt fails in Long CoT scenarios. To fill this gap, Li et al. [346] improve Vision RLLMs by enabling detailed, context-aware descriptions through an iterative self-refinement loop, allowing interactive reasoning for more accurate predictions without additional training. Dong et al. [126] incorporate multi-agent interaction during prompting, further scaling the reasoning length and achieving better accuracy. Furthermore, FaST [548] uses a switch adapter to select between Long CoT and direct answer modes, resulting in enhanced performance. (2) Multimodal Long CoT Imitation: Recent models such as LLaVA-CoT [713] and Virgo [133] employ data distillation to enable the imitation of Long CoT processes, addressing more complex problem-solving tasks [585]. Additionally, AtomThink [696] offers a Long CoT annotation engine that generates high-quality CoT annotations, mitigating the issue of insufficient visual mathematical data. Wei et al. [667] further extend Long CoT paradigms by incorporating more tokens during perception, improving geometric reasoning. (3) Reward Model-Based Multimodal Long CoT Exploration: Recent research employs reward or value models to enhance inference test-time scaling in both exploration and training phases. This includes model decoding [387, 52, 709] and RL training [696, 643, 815, 611, 228, 474, 558, 391], as well as the diffusion process [414, 774], all contributing to improved visual reasoning and comprehension.
The primary challenges in multimodal Long CoT are: (1) Incorporating Multimodal Reasonings: Enabling RLLMs to assist reasoning by generating [102, 182, 314] or grounding [682] visual content holds promise for improving complex spatial reasoning tasks, particularly when logic cannot be easily conveyed through text alone [314]. (2) Extending Longer Reasoning Processes: While current models focus on imitating Long CoT, there remains a lack of exploration into how multimodal inference test-time scaling can be achieved through methods like RL or MCTS [680, 241], presenting an interesting avenue for future research.
@@ 8.3 Agentic & Embodied Long CoT
Researchers have expanded Long CoT in interactive environments, significantly improving success rates in automated exploration tasks for agentic systems [185, 869, 815, 635, 637, 370] and embodied systems [803, 700, 844]. Current research primarily focuses on two approaches: (1) Tree-based Search Augmentation Early work by Hao et al. [185], Koh et al. [283] introduce tree search techniques to enhance agent exploration. Hu et al. [212] further propose planning sampling strategies to accelerate tree search processes. Additionally, Light et al. [356] develop a method to gather high-quality interactive feedback through self-play simulations with MCTS and LLM-based reflection, which helps acquire high-level strategic skills and guide low-level execution. (2) Environmental Interactivity Improvement A key feature of Agentic Systems is their understanding for the physical world [24, 278] and interaction with the environment [881, 141], making the enhancement of this aspect a critical focus [185]. Nie et al. [447] and Hu et al. [211] improve interactivity by incorporating memory history into the agent’s functions. (3) Multi-agent Cooperative Improvement Another key feature of agentic systems is that it can incorporate multiple agents to cooperative to solve a complex problem [903, 625, 482, 690, 901, 606, 763]. Christakopoulou et al. [108] introduce the Talker-Reasoner architecture, which separates the agent’s tasks into deep reasoning and rapid dialogue generation, providing a more effective interaction protocol. Lei et al. [305] introduce the Multi-Agent System for Conditional Mining (MACM) prompting method, which effectively addresses complex mathematical problems and exhibits robust generalization across diverse mathematical contexts.
@@ 8.4 Efficient Long CoT
The deep reasoning, exploration, and reflection of the Long CoT often lead to long outputs, which necessitate improved speedup techniques [749, 218, 156, 542, 244, 311]. Consequently, optimizing reasoning for faster reasoning with maximum accuracy has become a significant challenge for Long CoT [157]. Current research mainly focuses on two approaches: (1) Direct Compression and Shortening of Reasoning Chains: The most direct strategy is to consider direct compression and reducing the length of the reasoning chain while maintaining accuracy [104, 550, 22, 205]. Specifically, a series of work [572, 406, 59, 417, 444] encourage the generation of shorter reasoning processes, minimizing redundancy and enhancing efficiency [19, 720, 31, 442, 639, 393]. Additionally, researchers further introduce token budgets in prompts to control reasoning complexity, further improving efficiency [183, 808, 607, 243, 319, 5]. Building on these approaches, MARP [78] and DynaThink [454] allow LLMs to adapt reasoning speed based on task complexity, perplexity, or confidence, optimizing both efficiency and accuracy [171, 521, 906, 122, 116, 632, 269]. Moreover, Botta et al. [48] and Xia et al. [694] introduce a technique that enables LLMs to erase or skip some generated tokens, thereby compressing the reasoning length. More radically, Yu et al. [780] and Du et al. [130] propose distilling long reasoning paradigms into direct prediction models, reducing computational costs without sacrificing reasoning quality. (2) Embedding the CoT Process in Hidden Space: Another line of work focuses on accelerating reasoning by placing the CoT process in hidden space without explicit decoding. Specifically, Coconut [187], LaTRO [66], and SoftCoT [723] transfer reasoning into continuous latent space, promoting "continuous thinking" and enabling the model to maintain multiple alternative reasoning paths [830]. Similarly, Wang et al. [646] use "planning tokens" to enhance reasoning, performing the planning process in hidden space to save computational resources and improve inference performance.
@@ 8.5 Knowledge-Augmented Long CoT
The reasoning model significantly enhances reasoning capabilities, but it still lacks knowledge in specialized fields and timely new information [80, 138, 379, 537]. Thus, enriching reasoning with additional knowledge presents a key challenge for Long CoT [71, 64]. Current research focuses primarily on two approaches: (1) Retrieval-Augmented Generation: Retrieval-Augmented Generation (RAG) techniques enhance LLMs by integrating dynamic knowledge retrieval and document refinement [337, 647, 174, 254, 661]. Research has combined RAG with reasoning modules to improve performance on complex tasks [576, 259, 378, 685, 76]. O1 Embedder [727] optimizes multi-task retrieval and reasoning through synthetic data training. Furthermore, Stream of Search (SoS) [149] CoRAG [631] boost search accuracy and addresses unresolved issues by incorporating more natural reflection and exploration in RAG. (2) Model Knowledge Injection: An alternative approach involves integrating additional knowledge during SFT or RL [390]. Specifically, HuatuoGPT-o1 [71] utilize the R1-like paradigm to train LLMs by model-judged reward RL, which significantly improves the medical knowledge during reasoning [457, 229]. Huang et al. [233] and Wang et al. [615] optimize for injecting medical knowledge in Long CoT scenarios by SFT, which also achieve great performance. Further, Jiang et al. [256] introduce MCTS to synthesize data, achieving superior performance. This model merges verifiable medical knowledge with reinforcement learning techniques to enhance performance in complex, medical task settings.
@@
8.6 Safety and Stability for Long CoT
Despite the notable performance improvements brought about by Long CoT, Long CoT-augmented LLMs still encounter significant safety challenges. These include issues such as the generation of unstable outputs, exemplified by the tendency to memorize in-domain math questions instead of engaging in actual reasoning [726], and the production of unsafe outputs, such as misinformation and offensive content [890, 309, 889, 402, 20, 40, 39, 127, 275, 842]. Current research primarily addresses two key approaches: (1) Long CoT Attack Several studies show that Long CoT makes models more vulnerable to unexpected behavior [140] or unsafe outputs [288, 904, 719, 92, 18]. For instance, Arrieta et al. [21] identify that DeepSeek-R1 is prone to generating harmful content, including misinformation and offensive speech. Additionally, Kumar et al. [285] introduce the OverThink attack, which exploits false inference problems to induce overthinking in models, providing insights into potential defensive strategies. Further, Yao et al. [758] fool RLLMs chain of iterative chaos, for better jailbreaking. (2) Long CoT Safety Improvement Another major area of research focuses on enhancing safety [252, 900, 388] and reliability [565, 506, 598, 117, 89] through prompting [147] or training [459] techniques. Shen et al. [528] present Heima, which optimizes inference efficiency and robustness. Gallego [147] proposes dynamic security prompts during inference, while Cheng et al. [101] address hallucinations by guiding reasoning with a tree search algorithm. Zhao et al. [865] introduce a self-reflection framework to identify biases, and Wang et al. [620] propose Safety Reasoning with Guidelines (SRG) to defend against out-of-distribution attacks. Finally, Parmar and Govindarajulu [466] combine reinforcement learning (RL) and supervised fine-tuning (SFT) in a hybrid training approach to reduce harmful outputs and enhance DeepSeek-R1’s safety.